State the dimension of three rectangles with perimeter 32cm. Show your work to justify your answer
Answer by stanbon(75887)

State the dimension of three rectangles with perimeter 32cm. Show your work to justify your answer
-----------------
P = 2(L + W)
32 = 2(L+W)
L + W = 16
-----
The dimensions might be:
L = 14 and W = 2
L = 10 and W = 6
L = 5 and W = 11
or any other pair of numbers whose sum is 16.
-----
Cheers,
Stan H.
===================